[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed May 31, 2017
HSU Shift Log
nhc.tsbadmin at noaa.gov
Thu Jun 1 02:44:17 UTC 2017
HSU Shift Log for Wed May 31 2017 22:44:12 GMT-0400 (EDT)
Forecaster: D Brown
Technical Issues Summary:
ATCF issues: The EGRI track is not in the ATCF. Appears that this has
occurred much of the day. The 18z HWRF track only shows the initial
location of the system with no track thereafter, despite the model showing
a coherent vortex. Might be an issue.
Other Techincal issues:
Miscellaneous issues: Graphics at 03z were run immediately after the TCM
was issued. Apparently it needs the TCP as well. It plotted the new
forecast by showed the old time and previous watch area. Reran at 0240z.
Hopefully that will fix it! UW/CIMSS ADT page is listing the TD as an
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Ncep.nhc.nco_contacts